Solve for x. 2/5x = -10
Answer by tutorcecilia(2152)

2/5x = -10/1 [cross-multiply and solve for the x-term]
(2)(1)=(5x)(-10)
2=-50x
2/-50=x
-1/25=x
.
check by plugging (x=-1/25) back into the original equation and solve.